OpenStudy (anonymous):

what is the slope of the line that passes through the points (-2 5) and (1 4)

pooja195 (pooja195):

Use the slope formula \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] Plug in the numbers given in the coordinates like this: \[\LARGE \frac{4-5}{1--2}=slope\] two negative signs next to each other make a positive so.. \[\LARGE \frac{4-5}{1+2}=slope\] Do the math and you have your answer :-)

OpenStudy (anonymous):

Thank you!

pooja195 (pooja195):

yw ^_^ can you tell me what you got as the answer?

OpenStudy (anonymous):

-1/3

pooja195 (pooja195):

perfect :)